OpenStudy (anonymous):

Which values for x and y make the statement (x-5)(y + 6)=0 true? A. x=-5, y=6 B. x=5, y=-6 C. x=-5, y=-6 D.x=5, y=6

OpenStudy (mayankdevnani):

@chrisbush put all the values in option in given equation.If they satisfy the given equation, then that option is correct

OpenStudy (mayankdevnani):

can you get it what i am saying?

OpenStudy (anonymous):

No i dont understand because i was never taught this,im now in a corse recovery program for summer and i came across this awesome website (:

OpenStudy (mayankdevnani):

ok.. i will explain you

OpenStudy (mayankdevnani):

first take option A, where x=-5, y=6 then put these values in given equation The given equation is = \(\large (x-5)(y+6)=0\) so our new equation becomes, \[\large (-5-5)(6+6)=0\] now, solve for L.H.S .

OpenStudy (mayankdevnani):

@chrisbush

OpenStudy (anonymous):

Okay (: now how do i solve the rest of the new equation?

OpenStudy (mayankdevnani):

lol.. just like as option A . Only you need to do is that take different options like A,B and C. And solve for L.H.S and if any option's L.H.S is zero then that option is correct answer.

OpenStudy (mayankdevnani):

ok.. @chrisbush i hope you understand

OpenStudy (zzr0ck3r):

(-5-5)(6+6) = 0 is not true because (-5-5)(6+6) = -10*12 = -120 != 0 so this is not a solution, so not try the next

OpenStudy (anonymous):

Lol im so confused D; i dont know the answerrr

OpenStudy (mayankdevnani):

lolzz.... @chrisbush try again and again as @zzr0ck3r said

OpenStudy (zzr0ck3r):

notice if you put in x = 5 then the whole thing is 0 notice if you put in y = -6 then the whole things is 0 so pick the one/ones that have either x=5 or y=-6 or both

OpenStudy (zzr0ck3r):

note: 0*anything = 0

OpenStudy (anonymous):

so it would be b?

OpenStudy (mayankdevnani):

yes!!!

OpenStudy (anonymous):

lol no but thx i understand now

OpenStudy (mayankdevnani):

welcome:)

OpenStudy (zzr0ck3r):

I c I was reading it wrong, but you have it.:)

OpenStudy (anonymous):

thank you guys, very much appreciated (:

OpenStudy (mayankdevnani):

welcome:)